Grade 12Trigonometry

Prove that sin20°. sin40°. sin60°. sin80° = 3/16

Shalini  Rai
13 Years agoGrade 12

1 Answer

md salik
13 Years ago

=sin20 sin40 sin80 sin60

=1/4 sin60 sin 60          {sinx sin(60-x) sin(60+x)=1/4 sin 3x}

=1/4 *(sin 60)2

=1/4 *3/4

=3/16

proved